OpenStudy (anonymous):

Ann is twice as old as Jim was two years ago. The difference in their present ages is 7. Find both ages now.

OpenStudy (anonymous):

Let x be current age of Ann and y be the current age of Jim then, x=2(y-2) x-2y-4

OpenStudy (anonymous):

I think not enough information

ganeshie8 (ganeshie8):

@Rubie do you knw solving equations ?

ganeshie8 (ganeshie8):

x=2(y-2) -----------(1) x-y = 7 ------------(2)

ganeshie8 (ganeshie8):

you have two equations & two unknowns, can you solve them ?

OpenStudy (anonymous):

I'm not so sure abt the unknown equations though

OpenStudy (anonymous):

answer is 7&14resp